A proton is placed in uniform electric field that is directed downward and has a magnitude of 5.0 N/C. What is the magnitude and direction of the electric force experience by the proton?


1
Expert's answer
2022-02-16T15:05:10-0500

Given:

"E=5.0\:\rm N/C"

"e=1.6*10^{-19}\:\rm C"


The electric force on a charge

"F=eE=\rm 1.6*10^{-19}*5.0=8.0*10^{-19}\: N"

The force is directed downward.